# CommVQA: Situating Visual Question Answering in Communicative Contexts

### v1

@@ Abstract
Current visual question answering (VQA) models tend to be trained and evaluated on image-question pairs in isolation. However, the questions people ask are dependent on their informational needs and prior knowledge about the image content. To evaluate how situating images within naturalistic contexts shapes visual questions, we introduce CommVQA, a VQA dataset consisting of images, image descriptions, real-world communicative scenarios where the image might appear (e.g., a travel website), and follow-up questions and answers conditioned on the scenario. We show that CommVQA poses a challenge for current models. Providing contextual information to VQA models improves performance broadly, highlighting the relevance of situating systems within a communicative scenario.
@@ 1 Introduction
According to a pragmatic Bayesian view of communicative actions, people tend to ask questions that maximize the contextually relevant information gain based on their existing prior beliefs about the world Frank and Goodman (2012), following fundamental pragmatic principles Grice (1975). Based on these linguistic insights, we argue that prior VQA datasets largely do not consider two central communicative drives that limit their utility. The information people aim to obtain (and consequently the types of questions they ask) varies with (1) the person’s information needs based on their goals when encountering the image, and (2) the person’s prior knowledge of the image content. Thus we introduce CommVQA, a benchmark that treats VQA as an inherently communicative task.  [^1] All data and code will be made available at: https://github.com/nnaik39/commvqa.
Figure 1: Experimental design overview. First, images were crawled from Wikipedia and matched with plausible context scenarios. Descriptions for each image were initially elicited from GPT-4V and then edited by humans to ensure truthful and useful image descriptions. Then questions and answers were elicited from people, conditioned on the context and the image description, resulting in a minimum of three answers to each of the 1,215 unique visual questions. Instructions are simplified for illustration purposes (see Appendix A.1 for full prompts and task setups).
@@
We benchmark four state-of-the-art VQA models on CommVQA, and show that the situated nature of CommVQA poses a significant challenge for current state-of-the-art vision-language models (VLMs) and allows for analytical insights into model generation behavior due to a highly controlled data generation pipeline. Crucially, we find that models benefit from accessing the contextual information, further suggesting that these are important resources that shape VQA in communicative settings.
In summary, our main contributions are: (a) we introduce CommVQA, a benchmark consisting of images, contexts, descriptions, questions, and answers, where models must infer the details most relevant to the questioner’s goals, (b) we benchmark this dataset on current VQA approaches and explore whether models can be instructed to integrate this contextual condition, and (c) we analyze model performance on this task.
@@ 2.1 VQA Datasets
Image-Question-Answer Input
Most VQA datasets focus on image-question-answer triplets, which are constructed in isolation from the real-world contexts in which the images might appear Antol et al. (2015); Goyal et al. (2017); Hudson and Manning (2019); Marino et al. (2019). The VizWiz dataset Gurari et al. (2018) stands out among VQA datasets as it focuses on real-world accessibility. The images in VizWiz are elicited from BLV users uploaded a picture to a phone app with a question about their real-world environment. The goals behind the questions and images is specific to BLV people exploring their physical environment. CommVQA is complementary to VizWiz, since it investigates how contextually situating images online affects a model’s VQA task performance.
Image-Question-Answer-Context Input
Some VQA datasets have explored integrating supplementary information to the image. Visual Dialog Das et al. (2017) aims to create visual chatbots, which can answer a question based on an image and the prior dialogic context. Similarly to CommVQA, the questions in Visual Dialog can be similarly conceptualized as follow-up questions based on an image description. However, the images, questions, and answers are still decontextualized from a specific information goal beyond the questioner wanting to understand the image. CommVQA instead extends the focus on varying the questioner’s broader goals in understanding the image.
The ScienceQA Lu et al. (2022) and VQAOnline Chen et al. (2023) datasets were constructed by crawling online resources. ScienceQA Lu et al. (2022) consists of multimodal science understanding questions, where each multiple choice question and answer is associated with a lecture and explanation, functioning as the context. VQAOnline Chen et al. (2023) consists of images, questions, and context sourced from StackExchange. The question is the title and the context is the body of the StackExchange post, which constitutes supplementary information that’s useful for answering the question. The notion of context in these datasets is limited to texts that provide supplementary information, whereas we specifically focus on the effect of changing scenarios on the VQA task. Additionally, due to the controlled generation pipeline, CommVQA allows for insights on how models perform under varying contextual conditions.
@@ 3 The CommVQA Dataset
The CommVQA dataset was constructed in five main steps, as visualized in Figure 1. First, we sourced images from Wikipedia, and elicited both plausible scenarios and descriptions for the images from GPT-4V OpenAI (2023a). To ensure description quality, we conducted a human-subject study to edit the descriptions. We then elicited questions and answers for the dataset from US-based crowdworkers on Prolific. All human-subject studies were conducted with IRB approval.
@@ 3.1 Image-Scenario Matching
We extracted 150 images from Wikipedia pages of topics related to at least one of the scenario conditions. Since each image appears in two scenarios and each image-scenario pair is annotated with at minimum four questions and each question with three answers, this results in 3,642 unique question-answer pairs. We prioritized high coverage of individual datapoints over breadth since the focus of this work is on uncovering the diverse contextual effects on the VQA task.
@@
To assign plausible scenarios to the image, we instructed GPT-4V OpenAI (2023a) to rank the six scenarios in order of descending plausibility, using the following prompt, which was inspired by the framing in prior work Stangl et al. (2021): “Imagine you are a person browsing the Internet. Please rank the following scenarios in which you might encounter this image: 1) You are browsing a shopping website, with the goal of purchasing an item or experience…” (The full prompt to GPT4-V is included in Appendix A.1.) Using a human-subject pilot study, we established that GPT-4V made reasonable choices for contexts, which enabled us to easily scale. From the top-three scenarios for each image, two scenarios were selected in order to balance out the co-occurrence of contexts.
@@ 3.2 Image Description Elicitation
We elicited initial description drafts by prompting GPT4-V with the phrase: Describe this image in less than 150 characters to someone who cannot see it. The length constraint follows a commonly-issued guide on best-practices for accessibility alt text writing OSU (2024). We chose Wikipedia for sourcing images due to the copyright permissions and image variety.
@@
To ensure description quality, we conducted a human-subject study where participants edited the descriptions generated by GPT-4V OpenAI (2023a). These edits were intended to help balance for potential inaccuracies, possible misalignments with human description preferences, and the current design choices of GPT-4V. For instance, as of October 2023, the model refrains from explicitly identifying the people within an image OpenAI (2023b), even though proper names can be an important detail for a useful image description Bennett et al. (2021); MacLeod et al. (2017).
Each participant was shown six randomized trials with an image and description. They were instructed to edit the image description to correct any errors and make the it more useful to someone who cannot see it. Importantly, participants were not shown the context that each image was placed in to control for question variation across contexts. Participants could also choose to skip if no edits were needed. We recruited 44 participants and compensated them at the rate of \$13.50/hr.
We collected three description edits for each image description draft, and selected a random edited description for the final description, as long as the Jaccard distance between the original GPT-4V description and the edited description was less than $0.2$, to ensure that we weren’t selecting outliers.
@@ 3.3 Question Elicitation
To elicit visual questions, we recruited 139 participants, who were paid \$13.50/hr with an average completion time of seven minutes. In each trial, participants were given two pieces of information: an image description (e.g., “A group of people of various ages walking along a grassy path, with trees on one side.”), and a scenario for the image (e.g., Imagine you are browsing a Health website, with the goal of learning how to live a healthier lifestyle). Crucially, participants didn’t see the image to avoid priming for specific questions and simulate the visual inaccessibility of the image in the real-world scenario. Participants rated how likely the image would appear within the provided scenario, and were prompted to ask two questions they would like answered by someone who can see the image.
Through this setup, we elicited 1,569 questions in total, with at least four questions for each image-scenario pair.
@@ 3.4 Answer Elicitation
We elicited answers from 480 participants, who were paid \$13.50/hr, with an average completion time of seven and a half minutes. Participants were shown the image, question, context, and description. They were told that another user asked the question based on this image description, and asked to write an answer that would help the other person visualize the image. In the final dataset, each question was answered by at least three separate participants. We also included a checkbox for participants to indicate if a question was unanswerable. If a question was voted unanswerable by two or more annotators, we excluded the question from the dataset, resulting in 355 exclusions (22.6%).
@@ 4 Dataset Analysis
The final CommVQA dataset consists of 3,642 unique question-answer pairs, spanning across 300 unique image-scenario pairs. To better understand the challenges posed by this dataset, we now provide an analysis of CommVQA. Examples from the dataset, sampled to cover all scenarios, are given in Appendix Figure 6.
@@ 4.1 Analysis of Descriptions
In CommVQA, the descriptions are the basis for the VQA task and were collected in a two-stage process: first an automatic description generation by GPT-4V and then a human editing phase. The automatically generated descriptions had an average length of $99.36$ characters and in $50\%$ of trials, people didn’t make any edits to the descriptions. Additions had a minimal effect on length–when participants chose to make edits, they added extra information, resulting in an average length of $106.04$ characters.
@@ 4.2 Analysis of Questions
Figure 2: Heatmap displaying the accuracies that finetuned BERT achieved for classifying questions into each pair of scenarios. BERT achieved an accuracy of $91.5\%$ for distinguishing science magazines and shopping, yet BERT only achieved a performance of $86\%$ to distinguish questions between science magazines and news.
@@
If questions are context-sensitive, then a trained classifier should be able to predict the correct context from the question and achieve a performance reliably higher than random. To investigate whether this is true for CommVQA, we fine-tuned BERT Devlin et al. (2019) on the task of predicting whether each question appears within a certain context or not. We expect this task to be difficult even for a human, given that questions such as “What time of day is it?” might appear in any context. We split all the questions collected into 20 cross-validation folds, with an 80-10-10 train/test/val split, and fine-tuned for 200 epochs with LoRA Hu et al. (2021). If questions written in different contexts do not differ, we should achieve only a random performance at $16\%$. However, fine-tuned BERT achieved an accuracy of $57\%$, a significant improvement over random choice, showing that questions elicited within different contexts are distinguishable even from a language-only perspective.
While we find overall evidence that questions vary between all contexts, a classifier analysis further allows us to investigate which individual contexts have the least and most overlap in the questions asked. Figure 2 shows BERT’s performance when fine-tuned on distinguishing pairs of scenarios (e.g., shopping from science magazines), and indicates which scenarios are more easily distinguishable. Intuitively, certain contexts are more related than others–for instance, shopping and social media content is more likely to contain images of famous models, while shopping and science magazines are less likely to have overlap. Our analysis confirms this intuition and highlights where the translational accuracy of models might fail downstream.
We further inspect the question interrogatives for a potential source of between-context variation. The results of a Welch’s t-test demonstrated that, for example, compared to all other scenarios, “Who” questions are significantly more likely to appear in the social media scenario ($t(310)=2.86$, $p<0.001$), and “What” questions are significantly more likely to appear in the shopping scenario ($t(388)=2.86$, $p<0.005$) compared to any other scenario. These results indicate that different question types are more likely to be asked in certain scenarios, which carries implications that model evaluation should be contextual. A model that has poor performance on “Who” questions might seem competent in non-social media scenarios, but fail to generalize to social media due to the distinct nature of the user’s information needs.
@@ 4.3 Analysis of Answers
Firstly, with an average length of $9.8$ words, the answers in CommVQA are surprisingly long compared to prior datasets that used a similar online answer elicitation setup (e.g., 1.1 words for VQA-v2 Goyal et al. (2017) and 2.1 words for Visual Dialog Das et al. (2017)). Contrasting prior work, the instructions of CommVQA are framed as a communicative task, where answerers were asked to help someone else who cannot see the image. It’s therefore plausible that the longer responses are partially due to the participant’s wish to faithfully communicate with the questioner Grice (1975). In support of this hypothesis, we find that participants qualified many assumptions of theirs that were not directly supported by visual evidence with hedges and expressions of uncertainty, therefore lengthening the responses. In fact, the most common trigram within the dataset was “appear(s) to be” (93 occurrences), followed by “it looks like” (45 occurrences).
While we find that the scenario significantly shapes the content of the questions in CommVQA (see Section 4.2), it’s unclear to what extent we should expect the answers to be shaped by the scenario, as well. Since the questions are inherently affected by the context, it is difficult to establish a similar variation in the answers that isn’t already induced by the questions themselves. For example, we find that the rates at which named entities are mentioned largely varies between scenarios (see Appendix Figure 5), which is in line with the varied frequencies of “Who” questions. However, when we investigate the type of entities named in responses, we find additional effects of the scenario. For example, organizations are more frequently referenced in the travel scenario but individual people are more referenced in the science journal scenario (see Figure 3). This is borne out as a significant interaction between the proportion of mentioned organizations vs. people and the travel vs. science journal scenarios ($\beta=-2.22$, $SE=0.67$, $p<0.001$). This suggests that the external knowledge VQA models need to reference varies across scenarios and provides evidence that contextual considerations matter on the answer level.
Figure 3: Named entity recognition analysis showing significant variation within people’s answers between Travel and Science Journal scenarios. Organizations (e.g., companies, and institutions) were significantly more likely to be mentioned in the Travel over the Science scenario, and vice versa for people (e.g.., political figures, and celebrities). Asterisks indicate significant interaction. Error bars are 95% confidence intervals.

### v2

@@ Abstract
Current visual question answering (VQA) models tend to be trained and evaluated on image-question pairs in isolation. However, the questions people ask are dependent on their informational needs and prior knowledge about the image content. To evaluate how situating images within naturalistic contexts shapes visual questions, we introduce CommVQA, a VQA dataset consisting of images, image descriptions, real-world communicative scenarios where the image might appear (e.g., a travel website), and follow-up questions and answers conditioned on the scenario and description. CommVQA, which contains 1000 images and 8,949 question–answer pairs, poses a challenge for current models. Error analyses and a human-subjects study suggest that generated answers still contain high rates of hallucinations, fail to fittingly address unanswerable questions, and don’t suitably reflect contextual information. Overall, we show that access to contextual information is essential for solving CommVQA, leading to the highest performing VQA model and highlighting the relevance of situating systems within communicative scenarios.
@@ 1 Introduction
According to a pragmatic Bayesian view of communicative actions, people tend to ask questions that maximize the contextually relevant information gain based on their existing prior beliefs about the world Frank and Goodman (2012), following fundamental pragmatic principles Grice (1975). Based on these linguistic insights, we argue that prior VQA datasets largely do not consider two central communicative drives that limit their utility. The information people aim to obtain (and consequently the types of questions they ask) varies with (1) the person’s information needs based on their goals when encountering the image, and (2) the person’s prior knowledge of the image content. Thus we introduce CommVQA, a benchmark that treats VQA as an inherently communicative task. [^1] All data and code are available at: https://github.com/nnaik39/commvqa.
Figure 1: Overview of the CommVQA Dataset Construction. Images were sourced from Wikipedia and paired with relevant scenarios. The description were first generated by GPT-4V, then edited by humans. Other participants then provided questions and answers based on the scenario and description, resulting in at least three answers for each of the 2,983 unique visual questions. Simplified instructions are shown here; full details are in Appendix A.
@@
We benchmark four state-of-the-art VQA models on CommVQA, and show that the situated nature of CommVQA poses a significant challenge for current state-of-the-art vision-language models and allows for insights into model generation behavior due to a highly controlled data generation pipeline. We find that the most successful model needs to integrate contextual information, suggesting that context shapes VQA in communicative settings.
In summary, our main contributions are: (a) we introduce CommVQA, a benchmark consisting of images, contexts, descriptions, questions, and answers, where models must infer the details most relevant to the questioner’s goals, (b) we benchmark this dataset on current VQA approaches and explore whether models can be instructed to integrate context, and (c) we show via error analyses and a human-subjects experiment that even the best models generate false information at high rates.
@@ 2.1 VQA Datasets
Most VQA datasets focus on image-question-answer triplets, which are constructed in isolation from the real-world contexts in which the images might appear Antol et al. (2015); Goyal et al. (2017); Hudson and Manning (2019); Marino et al. (2019). The VizWiz dataset Gurari et al. (2018) stands out among VQA datasets as it focuses on real-world accessibility. The images in VizWiz are elicited from BLV users, who uploaded a picture to a phone app with a question about their real-world environment. The goals behind the questions and images is specific to BLV users exploring their physical environment. CommVQA is complementary to VizWiz, since it investigates how contextually situating images online affects a model’s VQA task performance.
Some VQA datasets have explored integrating supplementary information. Visual Dialog Das et al. (2017) aims to create visual chatbots, which can answer a question based on an image and the prior dialogic context. Similarly to CommVQA, these questions can be conceptualized as follow-up questions based on an image description. However, the images, questions, and answers are still decontextualized from a specific information goal beyond the questioner wanting to understand the image. CommVQA instead extends the focus on varying the questioner’s broader goals.
ScienceQA Lu et al. (2022) consists of multimodal science understanding questions, where each multiple choice question and answer is associated with a lecture and explanation, functioning as the context. VQAOnline Chen et al. (2023) consists of images, questions, and context sourced from StackExchange, where the context is the body of the post. The notion of context in these datasets is limited to texts that provide supplementary information, whereas we specifically focus on the effect of changing scenarios on the VQA task.
PromptCap Hu et al. (2022) similarly considers the notion of “context” for models generating answers in the VQA task. First, the model generates a visual description relevant to the question and then generates an answer solely relying on the description, conceptualized as “context.” In contrast to PromptCap, the contextual condition in CommVQA is strictly complementary to the image (see Section 3.3), therefore fundamentally changing the task and modeling demands.
@@ 3 The CommVQA Dataset
The CommVQA dataset was constructed in five main steps, as visualized in Figure 1. First, we sourced images from Wikipedia, and elicited both plausible scenarios and descriptions for the images from GPT-4V OpenAI et al. (2024). To ensure description quality, we conducted a human-subject study to edit the descriptions. We then elicited questions and answers for the dataset from US-based crowdworkers on Prolific. All human-subject studies were conducted with IRB approval.
@@ 3.1 Image-Scenario Matching
We extracted 1000 images from Wikipedia pages of topics related to at least one of the scenario conditions. We placed each image in two scenarios, which allows us to investigate how the scenario can induce variations in questions and answers across the same visual stimuli. Each image-scenario pair is annotated with on average 1.54 questions and each question with three answers, which results in 8,949 unique question-answer pairs. We prioritized high coverage of individual datapoints over breadth since the focus of this work is on uncovering the diverse contextual effects on the VQA task.
@@
To assign plausible scenarios to the image, we instructed GPT-4V OpenAI et al. (2024) to rank them in order of descending plausibility, guided by the task in Stangl et al. (2021). We validated the assignment on a subset of data with a human-subject study before using GPT-4V to scale (see Appendix A.2 for the prompt). From the top-three scenarios for each image, two scenarios were selected in order to balance out the co-occurrence of contexts.
@@ 3.2 Image Description Elicitation
We elicited initial description drafts by prompting GPT-4V with the phrase: Describe this image in less than 150 characters to someone who cannot see it. The length constraint follows a commonly-issued guide on best-practices for accessibility alt text writing OSU (2024). We chose Wikipedia for sourcing images due to the copyright permissions and image variety. All selected images were in the public domain.
@@
To ensure description quality, we conducted a human-subject study where participants edited the descriptions generated by GPT-4V OpenAI et al. (2024). These edits were intended to help balance for potential inaccuracies, possible misalignments with human description preferences, and the current design choices of GPT-4V. For instance, as of May 2024, the model refrains from explicitly identifying the people within an image OpenAI (2023), even though proper names can be an important detail for a useful image description Bennett et al. (2021); MacLeod et al. (2017).
Each participant was shown six randomized trials with an image and description. They were instructed to edit the image description to correct any errors and make it more useful to someone who cannot see it. Importantly, participants were not shown the context that each image was placed in to control for question variation across contexts. Participants could also choose to skip if no edits were needed. We recruited 369 participants and compensated them at the rate of \$13.50/hr.
We collected three description edits for each image description draft, and selected a random edited description for the final description.
@@ 3.3 Question Elicitation
To elicit visual questions, we recruited 619 participants, who were paid \$13.50/hr with an average completion time of seven minutes. In each trial, participants were given two pieces of information: an image description (e.g., “A group of people of various ages walking along a grassy path, with trees on one side.”), and a scenario for the image (e.g., Imagine you are browsing a Health website, with the goal of learning how to live a healthier lifestyle). Crucially, participants didn’t see the image to avoid priming for specific questions and simulate the visual inaccessibility of the image in the real-world scenario. They also rated how likely the image would appear within the provided scenario, and were prompted to ask two questions they would like answered by someone who can see the image.
In total, we elicited 2,983 questions, with 1.54 questions on average for each image-scenario pair. Based on a separate human-subject study (Section 5.2), we find that 80% of the questions require the image to answer, emphasizing the difficulty and inherent multimodal nature of CommVQA.
@@ 3.4 Answer Elicitation
We elicited answers from 870 participants, who were paid \$13.50/hr, with an average completion time of seven and a half minutes. Participants were shown the image, question, context, and description. They were told that another user asked the question based on this image description, and asked to write an answer that would help the other person visualize the image. Each question was answered by at least three separate participants.
We also included a checkbox for participants to indicate if a question was unanswerable. If a question was voted unanswerable by two or more annotators, we labeled the question as “unanswerable” within the dataset. In total, we collected 283 questions that were labeled unanswerable (9.5%). Within the naturalistic setting, where people will ask questions about an image they cannot see, it’s expected that people will ask unanswerable questions, and that models should have the capability to decline answering. The unanswerable portion of CommVQAcan help assess whether a model can abstain from answering instead of providing incorrect information MacLeod et al. (2017).
@@ 4 Dataset Analysis
The final CommVQA dataset consists of 8,949 unique question-answer pairs, spanning across 2000 unique image-scenario pairs. To better understand the challenges posed by this dataset, we now provide an analysis of CommVQA. Examples from the dataset, sampled to cover all scenarios, are given in Appendix Figure 5. We additionally conducted an analysis of a subset of the data and found similar patterns in model responses on this smaller dataset, suggesting our data is sufficiently large to obtain generalizable patterns. For more details, see Appendix C.
@@ 4.1 Analysis of Descriptions
In CommVQA, the descriptions are the basis for the VQA task and were collected in a two-stage process: first, automatic description generation by GPT-4V and then a human editing phase. The automatically generated descriptions had an average length of $63.663$ characters and in $42\%$ of trials, people didn’t make any edits to the descriptions. When editing, participants added extra information, increasing the average length to $97.29$ characters.
@@ 4.2 Analysis of Questions
Figure 2: Heatmap of BERT Classification Accuracy Across Scenario Pairs. When fine-tuned on different scenario pairs, BERT exhibits varying performance in its ability to distinguish between these scenarios. For instance, BERT achieved 94% accuracy when distinguishing between science magazines and shopping, but only $83\%$ accuracy for travel and social media.
@@
If questions are context-sensitive, then a trained classifier should be able to predict the correct scenario from the question and achieve a performance reliably higher than random. To investigate whether this is true for CommVQA, we fine-tuned BERT Devlin et al. (2019) on the task of predicting whether each question appears within a certain context or not. We expect this task to be difficult even for a human, given that questions such as “What time of day is it?” might appear in any scenario. We split all questions in the dataset into an 80-10-10 train/test/val split, and fine-tuned for 200 epochs with LoRA Hu et al. (2021). Fine-tuned BERT achieved an accuracy of $56\%$ on this task, a significant improvement over random choice (16%), showing that questions elicited within different contexts are inherently distinct.
While we find overall evidence that questions vary between all scenarios, a classifier analysis further allows us to investigate which individual scenarios have the least and most overlap in the questions asked. Figure 2 shows the BERT classifier’s performance when fine-tuned on distinguishing pairs of scenarios (e.g., shopping from science magazines), and indicates which scenarios are more easily distinguishable. Intuitively, certain scenarios are more related than others–for instance, shopping and social media content is more likely to contain images of famous models, while shopping and science magazines are less likely to have overlap. Our analysis confirms this intuition and highlights where models may fail to translate across scenarios.
We further inspect the question interrogatives for a potential source of between-scenario variation. The results of a Welch’s t-test demonstrated that, for example, compared to all other scenarios, “Who” questions are significantly more likely to appear in the social media scenario ($t(589.6)=3.85$, $p<0.001$), and “Where” questions are significantly more likely to appear in the travel scenario ($t(627.3)=3.58$, $p<0.0001$) compared to any other scenario. These results indicate that different question types are more likely to be asked in certain scenarios, which carries implications that model evaluation should be contextual. A model that has poor performance on “Who” questions might seem competent in non-social media scenarios, but fail to generalize to social media due to the distinct nature of the user’s information needs. However, question type and scenario are still distinct conceptually.
@@ 4.3 Analysis of Answers
Firstly, with an average length of $10.98$ words, the answers in CommVQA are surprisingly long compared to prior datasets that used a similar online answer elicitation setup (e.g., 1.1 words for VQA-v2 Goyal et al. (2017) and 2.1 words for Visual Dialog Das et al. (2017)). Contrasting prior work, the instructions of CommVQA are framed as a communicative task, where answerers were asked to help someone else who cannot see the image. It’s plausible that the longer responses are partially due to the participant’s wish to faithfully communicate with the questioner Grice (1975).
While we focus our analysis on the effects context directly has on the questions being asked (see Section 4.2), we also find that access to context becomes strictly necessary to answer those contextualized questions. Most strikingly, questions of the form “What else is in the image?” directly require the answerer to know what information users already have (see Figure 3 and Figure 6 for examples).